OpenStudy (anonymous):

How to factor q(x) =x^5-3x^4+3x^3+x^2-4x+2

OpenStudy (anonymous):

All I really need Is the first step!

OpenStudy (callisto):

Use factor theorem. q(1) = (1)^5 - 3(1)^4 + 3(1)^3 + (1)^2 -4(1) + 2 = 1 - 3 + 3 +1 - 4 + 2 = 0 So, (x-1) is a factor of q(x) Can you continue?

OpenStudy (anonymous):

Yeas thank you!

OpenStudy (anonymous):

Yes*